# Bayesian Optimization of Process Parameters of a Sensor-Based Sorting System using Gaussian Processes as Surrogate Models

### v1

@@ Abstract
Index Terms:
Sensor-Based Sorting, Process Optimization, Bayesian Optimization, Surrogate-Based Optimization, Gaussian Process RegressionI Introduction
@@ II-C Sorting Accuracy Determination
False Negatives (FN): Number of objects that belong to the reject class but were incorrectly sorted into the accept stream.
@@ III-B Bayesian Optimization using a Gaussian Process as Surrogate Model
III-C Strategy for Optimizing Process Parameter in Sensor-Based Sorting Systems Using Parallel Bayesian Optimization and a Combined Expected Improvement
@@
with weights . The parameter values for the highest combined EI were used as the settings for the subsequent sorting experiment.
Furthermore, an adjustment is made to consider the uncertainty of the results of the sorting experiments in the GPR model calculations. The variance of the measured values can be interpreted as noise and is integrated into the calculation by adding the individual variances as weights for each measured value to the diagonal of the kernel matrix as given by

### v3

@@ Abstract
I Introduction
@@ II-C Sorting Accuracy Determination
False Negatives (FN): Number of objects that belong to the accept class but were incorrectly sorted into the reject stream.
@@ III-B Bayesian Optimization using a Gaussian Process as Surrogate Model
III-C Weighted Multi-Objective Bayesian Optimization with Two Gaussian Processes and a Combined Expected Improvement
@@
with weights . The parameter values for the highest combined EI were used as the settings for the subsequent sorting experiment:
Furthermore, an adjustment is made to consider the uncertainty of the results of the sorting experiments in the GPR model calculations. We treat each measurement’s variance as observation noise. During GP training, we incorporate it by adding the scaled per-sample variances () to the kernel diagonal:
The factor controls how strongly the empirical noise estimates influence the fit.